OpenStudy (anonymous):

Related rates problem- Find the rate at which the volume of water in the pool is increasing at time t=8 hrs. How fast is the water level in the pool rising at t=8 hrs? Given..radius is 12ft, height of 4 feet; pool has 1000ft^3 water at t=0, During 0

OpenStudy (anonymous):

Our net rate of change of volume is given by \(\frac{dV}{dt}=p(t)-r(t)\); plug in \(t=8\) to determine our rate of volume rise. Then you can compute \(\frac{dh}{dt}\) after using implicit differentiation on the second equation:$$\frac{dV}{dt}=2\pi r\frac{dh}{dt}$$ ;-) Give us your table and I can show you.

OpenStudy (anonymous):

here ill just draw the table

OpenStudy (anonymous):